OpenStudy (anonymous):

What exponential function is the best fit for the data in the table? x f(x) 3 10 4 16 5 28

OpenStudy (ddcamp):

What tools are you allowed/supposed to use?

OpenStudy (anonymous):

tools? i'm not sure if we're supposed to use any tools but we basically just have to to figure out the equation

OpenStudy (anonymous):

the answer choices are a. f(x) = 3(2)^(x − 2) + 4 b. f(x) = 3(2)^(x − 2) − 4 c. f(x) = 1/3(2)^(x − 2) + 4 d. f(x) = 1/3(2)^(x − 2) − 4

OpenStudy (ddcamp):

Oh, choices, that makes things a bit easier.

OpenStudy (ddcamp):

Since you have choices, the first thing I would do is try each of them with 3, and see what works.

OpenStudy (tkhunny):

You will also have to define "best fit". What are the criteria for determining "best". How do you know one is better than another?